# MemInsight: Autonomous Memory Augmentation for LLM Agents

### v1

@@ 5.3 Event Summarization
For Claude-3-Sonnet, all metrics remain comparable, indicating that memory augmentations effectively capture the semantics and knowledge within dialogues at both the turn and session levels. This proves that the augmentations sufficiently enhance context representation for generating event summaries.
To further investigate how backbone LLMs impact augmentation quality, we employed Claude-3-Sonnet as opposed to Llama v3 for augmentation while still using Llama for event summarization. As presented in Table 7, Sonnet augmentations resulted in improved performance for all metrics, providing empirical evidence for the effectiveness and stability of Sonnet in augmentation.
@@ 6 Conclusion
This paper introduced MemInsight, an autonomous memory augmentation method that enhances LLM agents memory through attribute-based annotations. While maintaining comparable performance on standard metrics, MemInsight significantly improves LLM-based evaluation scores, highlighting its effectiveness in capturing semantics and boosting performance across tasks and datasets. Additionally, attribute-based filtering and embedding retrieval methods showed promising methods of utilizing the generated augmentations to improve the performance of various tasks. Priority augmentation enhancing similarity searches and retrieval. MemInsight also could be a complement to RAG models for customized retrievals, integrating LLM knowledge. Results confirm that attribute-based retrieval effectively enriches recommendation tasks, leading to more persuasive recommendations.
@@ 7 Limitations
While the proposed MemInsight model demonstrates significant potential in enhancing retrieval and contextual understanding, certain limitations must be acknowledged. MemInsight relies on the quality and granularity of annotations generated using LLMs, making it susceptible to issues such as hallucinations inherent to LLM outputs. Furthermore, although the current evaluation metrics provide valuable insights, they may not comprehensively capture all aspects of retrieval and generation quality, highlighting the need for the development of more robust and multidimensional evaluation frameworks.
@@ Appendix F Qualitative Analysis
Figure 9 illustrates the augmentations generated using different LLM models, including Claude-Sonnet, Llama, and Mistral for a dialogue turn from the LoCoMo dataset. As depicted in the figure, augmentations produced by Llama include hallucinations, generating information that does not exist. In contrast, Figure 10 presents the augmentations for the subsequent dialogue turn using the same models. Notably, Claude-Sonnet maintains consistency across both turns, suggesting its stable performance throughout all experiments. While Mistral model tend to be less stable as it included attributes that are not in the dialogue.

### v2

@@ 5.3 Event Summarization
For Claude-3-Sonnet, all metrics remain comparable, indicating that memory augmentations effectively capture the semantics within dialogues at both the turn and session levels. This proves that the augmentations sufficiently enhance context representation for generating event summaries. To further investigate how backbone LLMs impact augmentation quality, we employed Claude-3-Sonnet as opposed to Llama v3 for augmentation while still using Llama for event summarization. As presented in Table 7, Sonnet augmentations resulted in improved performance for all metrics, providing empirical evidence for the effectiveness and stability of Sonnet in augmentation. Additional experiments and detailed analysis are provided in Appendix E.4.
@@
Qualitative Analysis
To more rigorously assess the quality of the autonomously generated augmentations, we conduct a qualitative analysis of the annotations produced by Claude-3 Sonnet. Using the DeepEval hallucination metric Yang et al. (2024), we find that 99.14% of the annotations are grounded in the dialogue, demonstrating a high level of factual consistency. The remaining 0.86% primarily consist of abstract or generic attributes, rather than explicit inaccuracies. Additional experimental details and examples are provided in Appendix F.
@@ 6 Conclusion
This paper introduced MemInsight, an autonomous memory augmentation framework that enhances LLM agents’ memory through structured, attribute-based augmentations. While maintaining competitive performance on standard metrics, MemInsight achieves substantial improvements in LLM-based evaluation scores, demonstrating its effectiveness in capturing semantic relevance and improving performance across tasks and datasets. Experimental results show that both attribute-based filtering and embedding-based retrieval methods effectively leverage the generated augmentations. Priority-based augmentation, in particular, improves similarity search and retrieval accuracy. MemInsight also complements traditional RAG models by enabling customized, attribute-guided retrieval, enhancing the integration of memory with LLM reasoning. Moreover, in benchmark comparisons, MemInsight consistently outperforms baseline models in overall accuracy and delivers stronger performance in recommendation tasks, yielding more persuasive outputs. Qualitative analysis further confirms the high factual consistency of the generated annotations. These results highlight MemInsight’s potential as a scalable memory solution for LLM agents.
@@ 7 Limitations
While MemInsight demonstrates strong performance across multiple tasks and datasets, several limitations remain and highlight areas for future exploration. Although the model autonomously generates augmentations, it may occasionally produce abstract or overly generic annotations, especially in ambiguous dialogue contexts. While these are not factually incorrect, they may reduce retrieval specificity in tasks requiring fine-grained memory access. Additionally, MemInsight ’s performance is dependent on the capabilities of the underlying LLM used for attribute generation. Less capable or unaligned models may produce less consistent augmentations. We also acknowledge that our current implementation is limited to text-based interactions. Future work could extend MemInsight to support multimodal inputs, such as images or audio, enabling richer and more comprehensive contextual representations.
@@ E.2 Evaluation Framework
Model Llama v3 Mistral v1 Claude-3 Haiku Claude-3 Sonnet
Rel. Coh. Con. Rel. Coh. Con. Rel. Coh. Con. Rel. Coh. Con.
Baseline LLM Summary 2.23 2.66 2.63 3.34 3.77 4.11 3.97 4.33 3.79 3.27 3.64 2.78
MemInsight (TL) 1.60 2.17 1.95 2.53 2.49 2.38 3.98 4.37 3.66 3.09 3.27 2.77
MemInsight (SL) 1.80 2.62 3.67 4.09 4.38 4.19 3.94 4.31 3.69 3.08 3.39 2.68
MemInsight + Dialogues (TL) 2.41 2.79 3.01 4.30 4.53 4.60 4.24 4.43 4.16 3.25 3.43 2.86
MemInsight + Dialogues (SL) 2.01 2.70 3.86 4.04 4.48 4.34 3.95 4.33 3.71 3.02 3.37 2.73
Table 11: LLM-based evaluation scores for event summarization using relevance (Rel.), coherence (Coh.), and consistency (Con.) across different models and augmentation settings. Baseline summaries are generated using zero-shot prompting without memory augmentation. MemInsight is evaluated in both turn-level (TL) and session-level (SL) configurations, with and without access to dialogue context.
@@ E.3.1 Prompts
E.4 Additional Experiments
In this experiment, we include an additional baseline for event summarization: raw summaries generated directly by LLMs using zero-shot prompting, without any memory augmentation. This serves as a clear reference point to isolate the impact of MemInsight ’s augmentation strategy on summarization quality. Table 11 shows the results of this experiment. As illustrated, MemInsight consistently improves event summarization quality across models, with the best performance achieved when augmentations are integreted with dialogue context highlighting the value of fine-grained annotations and contextual grounding. Overall, the findings confirm that MemInsight enhances the factual and semantic quality of generated summaries.
@@ Appendix F Qualitative Analysis
Figure 9 illustrates the augmentations generated using different LLM models, including Claude-Sonnet, Llama, and Mistral for a dialogue turn from the LoCoMo dataset. As depicted in the figure, augmentations produced by Llama include hallucinations, generating information that does not exist. In contrast, Figure 10 presents the augmentations for the subsequent dialogue turn using the same models. Notably, Claude-Sonnet maintains consistency across both turns, suggesting its stable performance throughout all experiments. While Mistral model tend to be less stable as it included attributes that are not in the dialogue. A hallucination evaluation conducted using DeepEval yielded a score of 99.14%, indicating strong factual consistency. Table 12 presents examples of annotations with lower scores. While these annotations are more generic or abstract, they remain semantically aligned with the original input.
Input Augmentations Hall. Score
’Evan’: [[‘‘Evan’s son had an accident where he fell off his bike last Tuesday but is doing better now.", D20:3], [‘‘Evan is supportive and encouraging towards Sam, giving advice to believe in himself and take things one day at a time.", D20:9], [‘‘Evan is a painter who finished a contemporary figurative painting emphasizing emotion and introspection.", D20:15], [‘‘Evan had a painting published in an exhibition with the help of a close friend.", D20:17]], ’Sam’: [[‘‘Sam used to love hiking but hasn’t had the chance to do it recently.", D20:6], [‘‘Sam is struggling with weight and confidence issues, feeling like they lack motivation.", D20:8], [‘‘Sam acknowledges that trying new things can be difficult.", D20:12]] "evan":{"[event]":"<son’s accident>", "[emotion]":"<worry>", "[hobby]":"<hiking>", "[activity]":"<painting>"}, "sam":{"[emotion]":"<struggling>", "[issue]":"<weight>", "[emotion]":"<lack of confidence>", "[action]":"<trying new things>"}} 0.66
{’James’: [[‘‘James has a dog named Ned that he adopted and can’t imagine life without.", D21:3], [‘‘James is interested in creating a strategy game similar to Civilization.", D21:9], [‘‘James suggested meeting at Starbucks for coffee with John.", D21:13]], ’John’: [[‘‘John helps his younger siblings with programming and is proud of their progress", D21:2], [‘‘John is working on a coding project with his siblings involving a text-based adventure game.", D21:6], [‘‘John prefers light beers over dark beers when going out.", D21:16], [‘‘John agreed to meet James at McGee’s Pub after discussing different options.", D21:18]]} {"james":{"[emotion]":"<excited>", "[intent]":"<socializing>", "[topic]":"<dogs>", "[topic]":"<gaming>", "[topic]":"<starbucks>", "[topic]":"<pubMeeting>", "[activity]":"<coffee>", "[activity]":"<beer>"}, "john":{"[topic]":"<siblings>", "[topic]":"<programming>", "[activity]":"<adventure game>", "[emotion]":"<proud>", "[intent]":"<socializing>"}} 0.50
Table 12: MemInsight annotations that scored below 1% hallucination rate in the DeepEval hallucination evaluation.